Math 115. Homework 13 Due Thursday 2 December

Section 7.4 (Page 336): 1, 2, 3, 4, 5, 7 Section 7.5 (Page 340): 3, 4, 6 Additional problem #1 Prove the “stronger inequality” of Theorem 7.12 using the second assertion of Theorem 7.13 (instead of the method given in the book’s proof of Theorem 7.12). Possible erratum for Exercise 7.5.3 If you are using an older version of the textbook, please see the errata linked at the top of the course web page.